# Open Wireless Digital Twin: End-to-End 5G Mobility Emulation with OpenAirInterface and Sionna RT

### v3

@@ 3.1 RAY-TRACING SIMULATION RESULTS
In the Shin-nakano scenario, the radio propagation environment was complex owing to the presence of residential and low-rise buildings and narrow streets. Additionally, the 3D model provided by PLATEAU for the Shin-nakano area is at Level of Detail 1 (LOD1), resulting in relatively simple building shapes. In this mobility scenario, the vehicle started moving from a location far from the gNB at and stopped briefly at a traffic light between and . Upon receiving a green light, it makes a left turn and enters the street in front of the building where the gNB is installed. The vehicle continues eastward, passing in front of the gNB, briefly stops at another intersection, and then makes another left turn. It then proceeds to a shopping street heading north-northeast. This scenario consisted of 570 snapshots, corresponding to 57.0 s of mobility data. Using the CIR obtained every 100 ms, the power delay profile (PDP), calculated as the square of the complex amplitude, is shown in Fig. 4(b), and the phase delay profile is shown in Fig. 4(d). At approximately [s], as the vehicle enters the street in front of the gNB, the LoS path component appears. As the vehicle approaches the gNB, the signal strength increases and the delay decreases. When the vehicle passes the gNB and moves away, the signal strength weakens, the delay increases, and the LoS component exhibits a V-shaped pattern on the PDP. After the second left turn, not only does the LoS component diminish, but the number of reflection paths also drops significantly, resulting in approximately 30 dB of path gain attenuation. The change in the total path gain, obtained by summing all the complex amplitudes in the CIR, is shown in Fig. 4(c). The results clearly show a plateau period during the stop, a maximum value at approximately [s], and a sharp 30 dB drop at approximately [s]. Next, we discuss the phase behavior shown in Figs. 4(d) and (e). Fig. 4(d) indicates rapid phase fluctuations for each tap along the vertical delay-time axis. As shown in the magnified view in Fig. 4(e), the phase remained constant during the stop and then started fluctuating once the vehicle began moving. The phase exhibited a consistent trend along the LoS path trajectory observed in the PDP. However, it should be noted that this study does not account for time variations within the 100 ms intervals; consequently, the phase values show discontinuous transitions in each snapshot. This indicates that our study did not consider the Doppler effect.
@@ 3.2 OAI KPM RESULTS
Before applying the time variations of the channel induced by the mobility scenario, we performed a baseline measurement as a reference using the same experimental setup, as shown on the right side of Fig. 1 without applying any CIR. This means that, in the baseline measurement, in (4). The 100-second average values of the KPIs are summarized in Table 4.
@@
Figs. 6 and 7 summarize the time evolution of various KPIs output by the OAI during the execution of the Shin-nakano and Shibuya scenarios, respectively. The bands shown in lighter shades of the same color as each plot represent the upper and lower bounds observed when the corresponding scenario was executed ten times, indicating the range of the error margins.
@@
First, the variations in the RSRP obtained from the OAI and path gain from the Sionna RT, as shown in Figs. 6(a) and 7(a), exhibit a generally consistent trend. For example, in both scenarios, a plateau in the RSRP was observed during the red traffic light periods in both the DL and UL. However, the short-term fluctuations of approximately 20 dB and the long-term variations in the 18-44 second interval in the path gain of the Shin-nakano scenario are not fully reflected in the observed RSRP. This discrepancy is likely due to the KPI averaging process performed by the OAI gNB and the automatic gain control (AGC) function that may be implemented in a COTS UE. As a result, particularly in the Shin-nakano scenario, a level adjustment of approximately 20-30 dB might have occurred within the aforementioned time interval.
@@
Furthermore, the lower sections of these figures show that the MCS, BLER, and throughput fluctuate in response to time variations in the RSRP. In this section, we quantitatively examine the throughput values. The maximum bitrate [Mbps] of 5G NR can be estimated using the following formula, as specified in 3GPP TS 38.306 [58].
@@
Note that the notation is simplified by assuming the number of MIMO layers as 1, the number of aggregated component carriers as 1, and the scaling factor to be 1. Here, the modulation order , coding rate , and their product, the spectral efficiency are uniquely determined by the MCS index in TS 38.214 Table 5.1.3.1-2[59]. represents the number of physical resource blocks (PRBs) within , which is determined by the frequency range and numerology. is the average OFDM symbol duration in a subframe, and represents the control channel overhead. Some variables are determined by the parameters in Table 3 and remain unchanged throughout the scenario as , [µs], and is 0.14 for DL and 0.08 for UL. By substituting these specific values, the resulting equation is given in (8). The effective throughput for DL (UL) is defined as the actual received data rate, which is given in the case of TDD mode by
@@
where represents the block error rate defined by the ratio of the number of erroneous blocks to the total number of transmitted blocks, and indicates the time occupancy ratio of DL (UL) determined by the TDD slot pattern and special slot format. In the case of the TDD pattern shown in Table 3, the DL ratio is given by , while the UL ratio is given by , and thus, can be expressed as follows:
@@
Therefore, it can be concluded that the effective throughput can be calculated based on the spectral efficiency corresponding to the MCS index and . Accordingly, we focus on the variations in the MCS and depending on the mobility scenario. In the two scenarios analyzed, the maximum observed DL throughput was 150.0778 Mbps. At this point, the MCS and . Using (10), the calculated effective throughput is 155.2989 Mbps, showing a close match between measurement and calculation. On the other hand, the maximum observed UL throughput was 14.64214 Mbps. At this point, the MCS , and the . Using (11), the calculated effective throughput is 18.106756 Mbps. The discrepancy in UL throughput compared with DL throughput can be attributed to the presence of sounding reference signals (SRSs), which occupy a portion of the PRBs in UL transmission. Therefore, the KPIs measured by KPM xApps can be considered consistent with the theoretical expectations. The measured throughputs showed a maximum when the UE was closest to the gNB, the BLER exhibited an inverse trend to the throughput, and the MCS changed in a near-proportional relationship to the throughput, indicating highly reasonable results. When comparing Figs. 6(b) and (c), examining the difference between the DL and UL, it becomes apparent that UL variations are not as pronounced as those in the DL. The BLER is significantly higher, and to maintain UL communication quality, the MCS is automatically set lower to limit the modulation order, resulting in lower UL throughput. Potential causes of errors include (i) modulation and demodulation methods, (ii) modulation symbol period, (iii) characteristics of bandwidth limitation filters, (iv) channel characteristics, and (v) timing offsets due to cycle slips. These errors originate from both system factors and radio propagation issues. If we focus solely on the propagation aspects, the relevant factors are (iv) and (v). Because this study does not consider the velocity vector of the vehicle, the possibility of cycle slips caused by fast fading is eliminated. Consequently, inter-symbol interference (ISI) errors caused by multipath propagation with varying arrival times remain a primary concern. The Cyclic Prefix (CP) in CP-OFDM, used in 5G NR, is a technique designed to mitigate ISI. In our communication system parameters, the long CP assigned to the first OFDM symbol contains 132 samples corresponding to 2.86 µs, while subsequent short CPs contain 106 samples corresponding to 2.30 µs. If significant complex amplitudes are found in taps exceeding 132 samples within the 146 CIR taps used in this study, the throughput could be affected. In our environment, the real-time processing limit was taps, and no delay tap index exceeding 2.30 µs was detected. Thus, it is concluded that the ISI due to multipath propagation is not an issue in these scenarios, and the error observed in the UL is considered to originate from the system. As shown in Table 4, the UL BLER was already close to 10% during the baseline measurement, which may be related to the implementation of the received signal processing on the gNB.
@@
where is the RMS delay spread, is the guard-interval (CP) length, is the effective OFDM symbol duration (excluding the CP), is the fading (phase variation) period, is the maximum Doppler frequency, is the speed of light, and is the maximum UE speed. First, as discussed above, no multipath component with a delay exceeding the short CP length [µs] was observed; therefore, holds, and ISI is avoided. Second, the ratio (2.30 µs / 33.3 µs) follows the 5G NR standard, meaning and high spectral efficiency are maintained. Third, with the maximum vehicle speed [m/s], giving
@@
which is far longer than ; hence the phase coherence condition is also satisfied. Because all three inequalities are true, the Doppler shift [Hz] has a negligible effect on the link quality for CP-OFDM transmission under the parameter set of this study. At higher carrier frequencies (e.g., FR2) and high-speed mobility scenarios (e.g., trains), these conditions become far more stringent, and severe ISI or fast fading is expected. Because the shortest fading period is [ms], the CIR snapshot interval should, in principle, be no larger than . An efficient approach is to compute snapshots every approximately 20 ms (not exceeding several times ) and then use Sionna RT’s apply_doppler() function to interpolate the Doppler-affected CIR at a finer time resolution. Capturing the ms periodicity only requires interpolation at ms steps; therefore, the CIR data can be pre-generated offline. The remaining issue is to convolve these time-varying CIRs with the baseband signal in real-time. As the effective CIR interval shortens, the data loading overhead increases during real-time processing, necessitating additional acceleration techniques.
3.3 FUTURE PROSPECTS
@@
When discussing future intelligent transportation systems, it is necessary to consider methods for appropriately evaluating mobility scenarios in which the statistical properties of the channels change rapidly over time in dynamic radio environments. Although ray tracing achieves high-resolution results in static channel modeling, it becomes significantly more complex in dynamic scenarios involving high-speed moving vehicles, where the wide-sense stationarity (WSS) assumption is violated. Generally, for mobile entities in multipath environments, both frequency- and time-selective fading must be considered as doubly selective fading. However, realistically incorporating the effects of both types of fading into baseband signal processing and executing real-time E2E wireless communication is extremely challenging because of the computational load. To address this challenge, we aim to explore real-time processing, including support for Doppler effects, using acceleration technologies such as FPGAs and GPUs in future work.
@@
From a scaling perspective covering multiple gNB/UE nodes and MIMO emulation, we believe that the most promising approach is to offload the convolution operation to an FPGA board inserted in the cabled link between the gNB and UE. In this setup, the FPGA acts as a channel emulator that can process up to [the number of (#) RF ports on the board / 2] simultaneous channels (one TX/RX pair per channel). Accordingly, scalability is available as long as:
@@
Because convolution is well suited to the low-latency parallel architecture of an FPGA, we consider this hardware offload as an effective means to achieve scalable emulation.
@@
Beamforming for higher frequency bands, such as FR2, also necessitates tracking moving targets, which may lead to frequent handovers. Because a ray-tracing simulator can account for both the beam pattern of the gNB and the antenna pattern of the UE, it is possible to create FR2 mobility scenarios in which the beam dynamically tracks a moving UE. Once such a scenario is generated, the OWDT enables the development of a solution for managing handovers across the entire system.
@@
To achieve these objectives, the following extensions are required: integration with higher-precision 3D models and mobility scenarios, scalability in terms of the number of gNBs and UEs, utilization of acceleration technologies such as GPUs and FPGAs to enable real-time processing, and validation through comparison with experimental data to establish a truly representative digital twin.
4 CONCLUSION
@@
This study proposed the OWDT, an E2E 5G mobility emulation framework that can be deployed as an O-RAN ecosystem. By combining the OAI for protocol stack emulation and Sionna RT for ray-tracing-based RF modeling, the OWDT enables a high-fidelity performance evaluation in vehicular mobility scenarios. To achieve real-time emulation, we implemented a convolution method that applies pre-computed time-evolving CIRs to baseband IQ signals. This method efficiently reproduces multipath fading effects while ensuring real-time processing within the constraints of COTS hardware. The framework leverages near-RT RIC via FlexRIC to monitor KPIs, such as RSRP, MCS, BLER, and throughput, providing valuable insights into the behavior of wireless communication systems. Experimental validation demonstrated that the OWDT effectively replicates real-world channel conditions, supporting efficient network performance analysis and optimization with minimal physical deployment. This approach significantly reduces costs and accelerates the development of wireless communication systems. A key feature of this study is that all workflows and system architectures were built entirely using OSS and open data. The OWDT architecture we constructed serves as a reference architecture, and each component can be substituted with any proprietary product or alternative OSS implementation.
@@ ACKNOWLEDGMENT
T. Iye thanks Luis Pereira and Paulo Marques from Allbesmart LDA for the fruitful discussion on the OAI implementation. The authors used generative AI tools such as ChatGPT, Grammarly, and Paperpal to improve the grammar of the manuscript. After using these tools, they carefully reviewed and edited the content and took full responsibility for the final published manuscript.

### v4

@@ III-A RAY-TRACING SIMULATION RESULTS
In the Shin-nakano scenario, the radio propagation environment was complex owing to the presence of residential, low-rise buildings, and narrow streets. Additionally, the 3D model provided by PLATEAU for the Shin-nakano area is at LOD1, resulting in relatively simple building shapes. In this mobility scenario, the vehicle started moving from a location far from the gNB at and stopped briefly at a traffic light between and . Upon receiving a green light, it makes a left turn and enters the street in front of the building where the gNB is installed. The vehicle continues eastward, passing in front of the gNB, briefly stops at another intersection, and then makes another left turn. It then proceeds to a shopping street heading north-northeast. This scenario consisted of 570 snapshots, corresponding to 57.0 s of mobility data. Using the CIR obtained every 100 ms, the power delay profile (PDP), calculated as the square of the complex amplitude, is shown in Fig. 4(b), and the phase delay profile is shown in Fig. 4(d). At approximately [s], as the vehicle enters the street in front of the gNB, the LoS path component appears. As the vehicle approaches the gNB, the signal strength increases, and the delay decreases. When the vehicle passes the gNB and moves away, the signal strength weakens, the delay increases, and the LoS component exhibits a V-shaped pattern on the PDP. After the second left turn, not only does the LoS component diminish, but the number of reflection paths also drops significantly, resulting in approximately 30 dB of path gain attenuation. The change in the total path gain, obtained by summing all the complex amplitudes in the CIR, is shown in Fig. 4(c). The results clearly show a plateau period during the stop, a maximum value at approximately [s], and a sharp 30 dB drop at approximately [s]. Next, we discuss the phase behavior shown in Figs. 4(d) and (e). Fig. 4(d) indicates rapid phase fluctuations for each tap along the vertical delay time axis. As shown in the magnified view in Fig. 4(e), the phase remained constant during the stop and then started fluctuating once the vehicle began moving. The phase exhibited a consistent trend along the LoS path trajectory, as observed in the PDP. However, it should be noted that this study does not account for time variations within the 100 ms intervals; consequently, the phase values show discontinuous transitions in each snapshot. This indicates that our study did not consider the Doppler effect.
@@ III-B OAI KPM RESULTS
Before applying the time variations of the channel induced by the mobility scenario, we performed a baseline measurement as a reference using the same experimental setup, as shown in the right side of Fig. 1 without applying any CIR. This means that, in the baseline measurement, in (4). The 100-second average values of the KPIs are summarized in Table V.
@@
Figs. 6 and 7 summarize the time evolution of various KPIs output by the OAI during the execution of the Shin-nakano and Shibuya scenarios, respectively. The bands shown in lighter shades of the same color as each plot represent the upper and lower bounds observed when the corresponding scenario was executed ten times, indicating the range of error margins.
@@
First, the variations in the RSRP obtained from the OAI and path gain from the Sionna RT, as shown in Figs. 6(a) and 7(a), exhibit a generally consistent trend. For example, in both scenarios, a plateau in the RSRP was observed during the red traffic light periods for both the DL and UL. However, the short-term fluctuations of approximately 20 dB and the long-term variations in the 18-44 second interval in the path gain of the Shin-nakano scenario are not fully reflected in the observed RSRP. This discrepancy is likely due to the KPI averaging process performed by the OAI gNB and the automatic gain control (AGC) function that may be implemented in COTS UE. As a result, particularly in the Shin-nakano scenario, a level adjustment of approximately 20-30 dB might have occurred within the aforementioned time interval.
@@
Furthermore, the lower sections of these figures show that the MCS, BLER, and throughput fluctuate in response to time variations in the RSRP. In this section, we quantitatively examine the throughput. The maximum bitrate [Mbps] of 5G NR can be estimated using the following formula, as specified in 3GPP TS 38.306 [60].
@@
Note that the notation is simplified by assuming the number of MIMO layers as 1, the number of aggregated component carriers as 1, and the scaling factor to be 1. Here, the modulation order , coding rate , and their product, the spectral efficiency are uniquely determined by the MCS index in TS 38.214 Table 5.1.3.1-2[61]. represents the number of physical resource blocks (PRBs) within , which is determined by the frequency range and the numerology. is the average OFDM symbol duration in a subframe, and represents the control channel overhead. Some variables are determined by the parameters in Table IV and remain unchanged throughout the scenario as , [µs], and is 0.14 for DL and 0.08 for UL. By substituting these specific values, the resulting equation is given by (8). The effective throughput for DL (UL) is defined as the actual received data rate, which is given in the case of TDD mode by
@@
where represents the block error rate defined by the ratio of the number of erroneous blocks to the total number of transmitted blocks, and indicates the time occupancy ratio of DL (UL) determined by the TDD slot pattern and special slot format. In the case of the TDD pattern shown in Table IV, the DL ratio is given by , while the UL ratio is given by , and thus, can be expressed as follows:
@@
Therefore, it can be concluded that the effective throughput can be calculated based on the spectral efficiency corresponding to the MCS index and . Accordingly, we focus on the variations in the MCS and depending on the mobility scenario. In the two scenarios analyzed, the maximum observed DL throughput was 150.0778 Mbps. At this point, the MCS and . Using (10), the calculated effective throughput is 155.2989 Mbps, showing a close match between measurement and calculation. On the other hand, the maximum observed UL throughput was 14.64214 Mbps. At this point, the MCS , and the . Using (11), the calculated effective throughput is 18.106756 Mbps. The discrepancy in UL throughput compared with DL throughput can be attributed to the presence of sounding reference signals (SRSs), which occupy a portion of the PRBs in UL transmission. Therefore, the KPIs measured by KPM xApps can be considered consistent with the theoretical expectations. The measured throughputs showed a maximum when the UE was closest to the gNB, the BLER exhibited an inverse trend to the throughput, and the MCS changed in a near-proportional relationship to the throughput, indicating reasonable results. When comparing Figs. 6(b) and (c), examining the difference between the DL and UL, it becomes apparent that UL variations are not as pronounced as those in the DL. The BLER is significantly higher, and to maintain UL communication quality, the MCS is automatically set lower to limit the modulation order, resulting in lower UL throughput. The potential causes of errors include (i) modulation and demodulation methods, (ii) modulation symbol period, (iii) characteristics of bandwidth limitation filters, (iv) channel characteristics, and (v) timing offsets due to cycle slips. These errors originate from both system factors and radio propagation issues. If we focus solely on the propagation aspects, the relevant factors are (iv) and (v). Because this study does not consider the velocity vector of the vehicle, the possibility of cycle slips caused by fast fading is eliminated. Consequently, inter-symbol interference (ISI) errors caused by multipath propagation with varying arrival times remain a primary concern. The Cyclic Prefix (CP) in CP-OFDM, used in 5G NR, is a technique designed to mitigate ISI. In our communication system parameters, the long CP assigned to the first OFDM symbol contains 132 samples corresponding to 2.86 µs, while subsequent short CPs contain 106 samples corresponding to 2.30 µs. If significant complex amplitudes are found in taps exceeding 132 samples within the 146 CIR taps used in this study, the throughput could be affected. In our environment, the real-time processing limit was taps, and no delay tap index exceeding 2.30 µs was observed. Thus, it is concluded that the ISI due to multipath propagation is not an issue in these scenarios, and the error observed in the UL is considered to originate from the system. As shown in Table V, the UL BLER was already close to 10% during the baseline measurement, which may be related to the implementation of received signal processing on the gNB.
@@
where is the RMS delay spread, is the guard-interval (CP) length, is the effective OFDM symbol duration (excluding the CP), is the fading (phase variation) period, is the maximum Doppler frequency, is the speed of light, and is the maximum UE speed. First, as discussed above, no multipath component with a delay exceeding the short CP length [µs] was observed; therefore, holds, and the ISI is avoided. Second, the ratio (2.30 µs / 33.3 µs) follows the 5G NR standard, meaning and high spectral efficiency are maintained. Third, with the maximum vehicle speed [m/s], giving
@@
which is far longer than ; hence the phase coherence condition is also satisfied. Because all three inequalities are true, the Doppler shift [Hz] has a negligible effect on the link quality for CP-OFDM transmission under the parameters set in this study. At higher carrier frequencies (e.g., FR2) and high-speed mobility scenarios (e.g., trains), these conditions become far more stringent, and severe ISI or fast fading is expected. Because the shortest fading period is [ms], the CIR snapshot interval should, in principle, be no larger than . An efficient approach is to compute snapshots every approximately 20 ms (not exceeding several times ) and then use Sionna RT’s apply_doppler() function to interpolate the Doppler-affected CIR at a finer time resolution. Capturing the ms periodicity only requires interpolation at ms steps; therefore, the CIR data can be pre-generated offline. The remaining issue is to convolve these time-varying CIRs with the baseband signal in real time. As the effective CIR interval shortens, the data loading overhead increases during real-time processing, requiring additional acceleration techniques.
III-C FUTURE PROSPECTS
@@
When discussing future intelligent transportation systems, it is necessary to consider methods for appropriately evaluating mobility scenarios in which the statistical properties of the channels change rapidly over time in dynamic radio environments. Although ray tracing achieves high-resolution results in static channel modeling, it becomes significantly more complex in dynamic scenarios involving high-speed moving vehicles, where the wide-sense stationary (WSS) assumption is violated. Generally, frequency- and time-selective fading must be considered for mobile entities in multipath environments. However, realistically incorporating the effects of both types of fading into baseband signal processing and executing real-time E2E wireless communication are extremely challenging because of the computational load. To address this challenge, we aim to explore real-time processing, including support for Doppler effects, using acceleration technologies such as FPGAs and GPUs in future studies.
@@
From a scaling perspective covering multiple gNB/UE nodes and MIMO emulation, we believe that the most promising approach is to offload the convolution operation to an FPGA board inserted into the cabled link between the gNB and UE. In this setup, the FPGA acts as a channel emulator that can process up to [the number of (#) RF ports on the board / 2] simultaneous channels (one TX/RX pair per channel). Accordingly, scalability is available as long as:
@@
Because convolution is well suited to the low-latency parallel architecture of an FPGA, we consider hardware offloading as an effective means to achieve scalable emulation.
@@
Beamforming for higher frequency bands, such as FR2, also necessitates tracking moving targets, which may result in frequent handovers. Because a ray-tracing simulator can account for both the beam pattern of the gNB and the antenna pattern of the UE, it is possible to create FR2 mobility scenarios in which the beam dynamically tracks a moving UE. Once such a scenario is generated, the OWDT enables the development of a solution for managing handovers throughout the entire system.
@@
To achieve these objectives, the following extensions are required: integration with higher-precision 3D models and mobility scenarios, scalability in terms of the number of gNBs and UEs, utilization of acceleration technologies such as GPUs and FPGAs to support increased computational complexity, and validation through a comparison with experimental data to establish a truly representative digital twin.
IV CONCLUSION
@@
This study proposed the OWDT, an E2E 5G mobility emulation framework that can be deployed as an O-RAN ecosystem. By combining the OAI for protocol stack emulation and Sionna RT for ray-tracing-based RF modeling, the OWDT enables a high-fidelity performance evaluation in vehicular-mobility scenarios. To achieve real-time emulation, we implemented a convolution method that applies pre-computed time-evolving CIRs to baseband IQ signals. This method efficiently reproduces multipath fading effects while ensuring real-time processing within the constraints of COTS hardware. The framework leverages near-RT RIC via FlexRIC to monitor KPIs, such as RSRP, MCS, BLER, and throughput, providing valuable insights into the behavior of wireless communication systems. Experimental validation demonstrated that the OWDT effectively replicates real-world channel conditions, supporting efficient network performance analysis and optimization with minimal physical deployments. This approach significantly reduces costs and accelerates the development of wireless communication systems. A key feature of this study is that all workflows and system architectures were built entirely using OSS and open data. The OWDT architecture we constructed serves as a reference architecture, and each component can be substituted with any proprietary product or alternative OSS implementation.
@@ ACKNOWLEDGMENT
T. Iye thanks Luis Pereira and Paulo Marques from Allbesmart LDA for the fruitful discussion on the OAI implementation. The authors used generative AI tools, such as ChatGPT, Grammarly, and Paperpal, to improve the grammar of the manuscript. After using these tools, they carefully reviewed and edited the content and took full responsibility for the final published manuscript.